OpenStudy (anonymous):

Use substitution to solve the system. Use the substitution method.. 5x-3y=21 y=2x-8 What does x and y equal

OpenStudy (surry99):

so 5x -3 (2x-8) = 21 5x - 6x + 24 =21 -x = -3 x = 3 and y = 2(3) -8 = -2

OpenStudy (anonymous):

OMG thank you so much I have two more problems like these.. Can you answer them for me

OpenStudy (surry99):

sure

OpenStudy (anonymous):

3y-11=x 3x+5y=9

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

3y-11=x 3x+5y=9 3(3y-11)+5y=9 9y-33+5y=9 14y=9+33 14y=42 y=42/14=3 x=3y-11=3*3-11=9-11=-2

OpenStudy (anonymous):

y=3 and x=-2

OpenStudy (anonymous):

omg thank you so much

OpenStudy (anonymous):

I have one more, will you help me

OpenStudy (anonymous):

sure ^_^

OpenStudy (anonymous):

-2x+y=1 5x+4y=17

OpenStudy (anonymous):

please help @shahd-owies

OpenStudy (surry99):

rearrange the first to give: y = 1 + 2x then 5x + 4(1 + 2x) = 17 5x + 4 + 8x = 17 13x = 13 x =1 y = 1 + 2x = 1 + 2(1) = 3

OpenStudy (anonymous):

-2x+y=1 y=1+2x 5x+4y=17 5x+4(1+2x)=17 5x+4+8x=17 13x=13 so x=1 y=1+2x=1+2=3 y=3

OpenStudy (anonymous):

thanks so much you guys